A cylinder contains 12.4 L of air at a total pressure of 41.4 psi and a temperature of 35 ∘C.  How many moles of gas does the cylinder contain? (Hint: You must convert each quantity into the correct units (L, atm, mol, and K) before substituting into the ideal gas law.)
Answer #1

P = 41.4 psi = 41.4×0.068 atm = 2.82 atm

(1psi = 0.068 atm)

V = 12.4 L

T = 35°C = 35 + 273.15 K = 308.15 K

R = 0.082 Latm/mol.K

n = PV/RT = 2.82×12.4/(0.082×308.15) mol

= 1.38 mol